The acts in a talent competition consist of 4 instrumentalists, 10 singers, and 6 dancers. If the acts are ordered randomly, what is the probability that a dancer performs first?


1
Expert's answer
2021-03-01T15:59:59-0500

There are 4 instrumentalist, 10 singers and 6 dancers.


Total number of participant = 20


Since There are random selection each talent type can occur with probability

"P(1)=P(2)=P(3)=\dfrac{1}{3}"


Probability of selecting the instrumentalist

"P(I)=\dfrac{4}{20}=\dfrac{1}{5}"



Probability of selecting the Singers

"P(S)=\dfrac{10}{20}=\dfrac{1}{2}"



Probability of selecting the Dancers

"P(D)=\dfrac{6}{20}=\dfrac{3}{10}"



Probability that Dancer performs first="\dfrac{P(1).P(D)}{P(1).P(D)+P(2).P(I)+P(3).P(D)}"


"=\dfrac{\dfrac{1}{3}.\dfrac{3}{10}}{\dfrac{1}{3}.\dfrac{3}{10}+\dfrac{1}{3}.\dfrac{1}{5}+\dfrac{1}{3}.\dfrac{1}{2}}"


"=\dfrac{3}{10}=0
.3" .